OpenStudy (anonymous):

Which of the following equations correctly represents a circle centered at the origin with a radius of 5? A. x2 + y2 = 25 B. x2 + y2 = 5 C. (x - 5)2 + y2 = 25 D. (x - 5)2 + (y - 5)2 =5

OpenStudy (anonymous):

are you going to post the choices

OpenStudy (anonymous):

The general equation for a circle is: \[(x-h)^2+(y-k)^2=r^2\] Where h and k represent the coordinates of the center. If it's the origin, it simplifies to: \[x^2+y^2=r^2\]

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

remeber it's equal to r^2, so if r=5...

OpenStudy (anonymous):

it's not d.

OpenStudy (anonymous):

A